Intro Your grandmother left you $40,000 that youll receive when you turn 40 in 21 years. Since youd rather have some money
Answer #1

This can solved by using simple Future Value Formula:

let present value be x.

Time Horizon is 21 years.

A

Future Value = Present Value(1+r)t  

40000 = x (1+0.06)21

x = 40000/ 3.39956

x = $ 11766.23

B

40000 = x (1+0.08)21

x     = 40000/5.0338

x =$ 7946.28

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you like the answer, do upvote. Help us help you.